OpenStudy (emsxx3_):

Line BC contains points B (4, −5) and C (3, 2). Line DE contains points D (2, 0) and E (9, 1). Lines BC and DE are A. parallel B. perpendicular C.neither

OpenStudy (luffingsails):

Figure out the slope of the two lines and then compare them. That will help you answer the question. Can you start with the slope?

OpenStudy (emsxx3_):

idk the slope @luffingsails

OpenStudy (luffingsails):

The slope is defined as the rise/run. Or the difference in y values divided by the difference in x values. So, the slope of line BC = (-5-2) / (4-3) = -7/1 = -7. Can you try the slope of line DE now?

OpenStudy (emsxx3_):

so it would be (9 -2) / (0 -1) = 7/-1 = -7

OpenStudy (luffingsails):

close, remember that the differences in y values goes in the numerator. So, (1-0)/(9-2) = 1/7

OpenStudy (emsxx3_):

oh okay

OpenStudy (luffingsails):

So, we are looking at the two slopes -7 and 1/7 and comparing them. If they had the same slope, they would be parallel. But they don't, so they aren't.

OpenStudy (luffingsails):

There is a case where if one slope is the negative reciprocal of the other then they are perpendicular. That is the case here. If I started with -7, took the reciprocal (-1/7), and then multiplied by negative 1 (which gives us 1/7)... that proves they are perpendicular.

OpenStudy (emsxx3_):

so its c?

OpenStudy (luffingsails):

Sorry, was typing my answer. No, the two lines are perpendicular. So, B.

OpenStudy (emsxx3_):

its okay

OpenStudy (emsxx3_):

and thank you

OpenStudy (luffingsails):

Welcome.